OpenStudy (aaronandyson):

if x and y are two real unequal quantities then prove that cos theta = 1 + 1/x is impossible

OpenStudy (aaronandyson):

\[\cos \theta = 1 + \frac{ 1 }{ x }\]

OpenStudy (loser66):

What is the role of y here?

OpenStudy (aaronandyson):

@mayankdevnani

OpenStudy (aaronandyson):

something is wrong with my account..read th question again

OpenStudy (mathmate):

@AaronAndyson hints: 1. find the range of cosine(x). 2. prove that 1+1/x is outside of range of cosine(x)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

there is a problem with the variables here

OpenStudy (misty1212):

if for example \(x=-2\) then \[1+\frac{1}{x}=\frac{1}{2}\]which is not outside the range you have \(x,y,\theta\) too many variables in the question

OpenStudy (mathmate):

@AaronAndyson Quite true @misty1212! Is the question exactly what you posted, or are there minor differences? is x a positive number, or is cos(x)=x+1/x ?? (this is what I was working with, sorry). By the way, where does y come in?